# Better Rigs, Not Bigger Networks: A Body Model Ablation for Gaussian Avatars

### 01 Apr 2026 v1

@@ 2.2 Parametric Body Models
Momentum Human Rig offers a substantially richer mesh representation with 127 supplemental twist joints, 18,439 vertices, non-linear pose corrections, decoupled skeleton-mesh architectures and a partitioned identity space [2]. Its twist joints directly address the candy-wrapper artifacts that SMPL-based pipelines must learn to compensate for, and its denser mesh provides finer-grained Gaussian attachment. In order to help estimate MHR parameters, SAM-3D-Body was also released which enables fitting of a mesh from a single image [21]. Despite MHR’s representational advantages, no prior avatar method, to our knowledge, has explored replacing SMPL with a higher-capacity body model.
@@ 4.2 Main Results
The narrower margins on SSIM and LPIPS are expected as both metrics are more sensitive to perceptual high-frequency texture quality, precisely the phenomena that color MLPs and pose-conditioned decoders in prior methods target. We deliberately disable these components to isolate the body model’s contribution and provide a floor for MHR’s performance. The remaining gap suggests that lightweight learned corrections on top of MHR would likely close it, a direction we leave to future work.
@@ 4.4 Disentangling Pose Quality and Mesh Capacity
Our main results conflate two factors: SAM-3D-Body may produce more accurate poses than the Anim-NeRF fits used by prior methods, and MHR’s mesh may better support Gaussian attachment than SMPL-X. Figure 2 illustrates the pose estimation gap. The SAM-3D-Body fit (right) captures head orientation, eye gaze, and foot placement more faithfully than the Anim-NeRF fit (left), and its mesh contours follow the underlying human shape rather than the clothing silhouette. Table 3 isolates each factor on the two PeopleSnapshot subjects for which Anim-NeRF’s optimized SMPL parameters are publicly available via links provided by the Anim-NeRF team. Averaged across both subjects, upgrading pose estimation alone accounts for +2.50 PSNR, +0.0123 SSIM, and -0.0070 LPIPS, while upgrading the rendering mesh contributes a further +2.41 PSNR, +0.0080 SSIM, and -0.0080 LPIPS, confirming that both factors provide independent, roughly comparable gains.

### 03 Apr 2026 v2

@@ 2.2 Parametric Body Models
Momentum Human Rig offers a substantially richer mesh representation with 127 joints, 18,439 vertices, non-linear pose corrections, decoupled skeleton-mesh architectures and a partitioned identity space [2]. Its twist joints directly address the candy-wrapper artifacts that SMPL-based pipelines must learn to compensate for, and its denser mesh provides finer-grained Gaussian attachment. In order to help estimate MHR parameters, SAM-3D-Body was also released which enables fitting of a mesh from a single image [21]. Despite MHR’s representational advantages, no prior avatar method, to our knowledge, has explored replacing SMPL with a higher-capacity body model.
@@ 4.2 Main Results
The narrower margins on SSIM and LPIPS are expected as both metrics are more sensitive to perceptual high-frequency texture quality, precisely the phenomena that color MLPs and pose-conditioned decoders in prior methods target. We deliberately disable these components to isolate the body model’s contribution and provide a floor for MHR’s performance. The remaining gap suggests that lightweight learned corrections on top of MHR would likely be useful, a direction we leave to future work.
@@ 4.4 Disentangling Pose Quality and Mesh Capacity
Our main results conflate two factors: SAM-3D-Body may produce more accurate poses than the Anim-NeRF fits used by prior methods, and MHR’s mesh may better support Gaussian attachment than SMPL-X. Figure 2 illustrates the pose estimation gap. The SAM-3D-Body fit (right) captures hand orientation, head shape, and foot placement more faithfully than the Anim-NeRF fit (left), and its mesh contours follow the underlying human shape rather than the clothing silhouette. Table 3 isolates each factor on the two PeopleSnapshot subjects for which Anim-NeRF’s optimized SMPL parameters are publicly available via links provided by the Anim-NeRF team. Averaged across both subjects, upgrading pose estimation alone accounts for +2.50 PSNR, +0.0123 SSIM, and -0.0070 LPIPS, while upgrading the rendering mesh contributes a further +2.41 PSNR, +0.0080 SSIM, and -0.0080 LPIPS, confirming that both factors provide independent, roughly comparable gains.